OpenStudy (qwerd722):

Integrate: (tan x/3 + cot x/3)^2 dx

oregonduck (oregonduck):

Welcome To OpenStudy So you know some people give direct answers but if you decide to help people lead them to the answer. Do not expect people to come and give you the answer. Please be able to work it out with your helper. If you can not work with your helper no one will help you. Please make sure you read OpenStudy's Code Of Conduct Once again WELCOME TO OPENSTUDY!

OpenStudy (kainui):

This is quite a fun problem, at least the way in which I figured it out: First, I just squared it: \[\int (\tan \tfrac{x}{3} + \cot \tfrac{x}{3})^2dx\] \[\int \tan^2 \tfrac{x}{3} +2+ \cot^2 \tfrac{x}{3} dx\] notice that the middle term there comes from \(2 \tan \tfrac{x}{3} \cot \tfrac{x}{3} = 2\). Now from here I recognized I could plug in the Pythagorean identity in two different ways: \[\tan^2 \tfrac{x}{3} + 1 = \sec^2 \tfrac{x}{3}\] for example with the other one for cotangent to get: \[\int \sec^2 \tfrac{x}{3} + \csc^2 \tfrac{x}{3} dx\] And now you should be able to figure out the rest.